OpenStudy (anonymous):

Consider:2(-3i+4)-i(4-i2) Simplify the expression and then determine a and b that produce an equivalent expression of the form a+bi. a. A=8,b=11 b. A=8,b=-11 c. A=7,b=-10 d. A=7,b=9

OpenStudy (michele_laino):

first step: since \(i^2=-1\), then we can rewrite such expression as below: \[\Large \begin{gathered} 2\left( { - 3i + 4} \right) - i\left( {4 - {i^2}} \right) = \hfill \\ \hfill \\ = - 6i + 8 - i\left( {4 + 1} \right) = ...? \hfill \\ \end{gathered} \] please continue

OpenStudy (anonymous):

what?